The sum of the parent's age is twice the sum of their children's ages. Five years ago, the sum of parent's ages is four times the sum of children ages. In fifteen years, the sum of the parent's ages will be equal to the sum of their children's ages. How many children were in the family?
Answer by mananth(16949)

let the children in the family be x
their sum of ages = y
let the parents sum of ages be z
z=2y
.
FIVE YEARS AGO
z-10 = 4(y-5x)
z-10=4y-20x
but z=2y
plug the value of z
2y-10=4y-20x
2y-20x=-10------------ A
.
.
15 YEARS LATER
z+30 = y+15x
2y+30=y+15x
y-15x=-30-------------B
2(y-15x)-(2y-20x)= 2*-30 - 2*-10
2y-30x-2y+20x=-60+20
-10x=-40
x=4
there were 4 children